OpenStudy (anonymous):

evaluate the function for the given value g(-2)for g(x)= -2x^3-1

OpenStudy (anonymous):

Just substitute x=-2 in the function.

OpenStudy (anonymous):

i tried and dont get an accurate answer :(

OpenStudy (anonymous):

I think your issue may be that you aren't using parentheses. So try -2 (-2)^3 - 1 which will give you -2*-8 -1 = 16-1=15 hope this helps!!

OpenStudy (anonymous):

THANK YOU